A sports psychologist tested whether the number of team wins in the previous season can predict the number of wins in the following season in a sample of 35 teams. Complete the regression table for this hypothetical study. (Round your value for Fobt to two decimal places.)

Source of
Variation
SS df MS Fobt
Regression 49
Residual (error) 6
Total
Answer #1

df(Regression) = Number of independent variable = 1

df(Total) = n - 1 = 35 - 1 = 34

df(Error) = 34 - 1 = 33

MS(Regression) = 49/1 = 49

SS(Residual) = 34*6 = 204

SS(Total) = 49 + 204 = 253

Fobt = 49/6 = 8.17

Hence,

The complete table will be:

Source of
Variation
SS df MS Fobt
Regression 49 1 49 8.17
Residual (error) 204 33 6
Total 253 34